OpenStudy (anonymous):

how do you find the area between 2 curves?

OpenStudy (anonymous):

what grade you in?

OpenStudy (amistre64):

take the integral of the top and subtract the integral of the bottom.

OpenStudy (amistre64):

In essense, area of the bigger (-) area of the smaller gives you area left between curves.

OpenStudy (anonymous):

hi amistre64, say if these are the curves y=x^2+5x & y=3-x^2 between x=-2 & x=0...then?

OpenStudy (amistre64):

f(x) = x^2 +5x and g(x) = 3 -x^2 (S) f(x) dx = (1/3)x^3 + (5/2)(x^2) + x; F(0) - F(-2) = area under f(x). (S) g(x) dx = 3x - (1/3)x^3; G(0) - G(-2) = area under g(x) Subtract the area to determine area between the curves :)

OpenStudy (amistre64):

You can forgo the F(0) and G(0), they are 0 to begin with :)

OpenStudy (anonymous):

(S) f(x) dx = (1/3)x^3 + (5/2)(x^2) + x <<<<<<<why did u add this 'x' ?

OpenStudy (anonymous):

Hey since we are in the same level of math, can someone help me with converge and diverge?

OpenStudy (amistre64):

cause im old and forgetful :) forget that little "x"....

OpenStudy (anonymous):

:) thnx..can u temme how to find which is the upper curve & which one is the lower one? any hints/tricks?

OpenStudy (amistre64):

If you know the graphs, its easier; but to tell the truth I dont think it matters much. Let me see.... if were going with a box. And one box is 50 and the other is 20, we would want to find the difference between them. 50-20 = 30; and 20-50=-30. Intuition says that if the area is negative, it is just a simple change of sign to make it right :)

OpenStudy (amistre64):

Does it make sense to you?

OpenStudy (anonymous):

yeah..of course!!

OpenStudy (anonymous):

just see the "ask a question " pane

OpenStudy (anonymous):

hey , can't we ask more than one question ? i am unable to click on "Ask" there...!

OpenStudy (anonymous):

my question is ,whats the difference between total derivative & total differentiation ?

OpenStudy (amistre64):

I have never been able to click on it, it just freezes kindof. And these posts are getting sticky too. Click "post" and have to come back to the "room" to type something else.... What I wouldnt give for a pencil and paper these days :)

OpenStudy (anonymous):

yeah its the same here...it's like stuck!!

OpenStudy (amistre64):

I dunno about the "total" part. But a derivative is when you use the symbol (dy/dx) or something like it. The derivative (dy/dx) cannot be a number, or seperated. Differentiation on the other hand, is using the (dy/dx) as if it WAS a fraction and seperating it. For examples: dy/dx = x^2 is a derivative. dy = x^2 dx is a deriviation. Why there is a difference, i dont think ill ever know.

OpenStudy (anonymous):

example for differentiation?

OpenStudy (anonymous):

i need help in my mathematical literacy homework

OpenStudy (amistre64):

There are alot of "d" words to keep track of in Calculus, youd think theyd have come up with better names... dy = x^2 dx is an example of differentation. And: It is the differentation that you work the integral on. (S) dy = (S) x^2 dx y = (1/3) x^3

OpenStudy (amistre64):

mathmatical literacy? what would that consist of?

OpenStudy (amistre64):

I might be confusing some issues there, so its always best to get a second opinion :)

OpenStudy (anonymous):

algbra

OpenStudy (anonymous):

amistre64, are you a professor somewhere?

OpenStudy (amistre64):

kabela; are you able yo use the "ask a question" box?

OpenStudy (amistre64):

I am getting a Masters in Math so I can be a college Math teacher and warp young fragile minds :)

OpenStudy (anonymous):

ohhh...we are in the same level then :) great!

OpenStudy (amistre64):

im taking "Trig" right now in college, and I am sooooo bored with it. I taught myself trig 3 semesters ago when they had me in intermediate algebra. So now I am teaching myself the Calculus stuff...

OpenStudy (anonymous):

ohh...grt...then u can always clear my doubts..as i have calculus as my subject now...:) can i have your email id if possible?

OpenStudy (amistre64):

just add "@gmail.com" to my screenname.

OpenStudy (anonymous):

okay..thanks!!!...is ur mailbox open right now? i am dropping a mail to ur inbox

OpenStudy (amistre64):

I can open it if need be. I remember the password and everything :)

OpenStudy (anonymous):

just chk ur inbox..n respond so that i may know ive sent to the correct id

OpenStudy (anonymous):

did u get it?

OpenStudy (amistre64):

got it....